# fPLSA: Learning Semantic Structures in Document Collections Using Foundation Models

### v1

@@ Abstract
Humans have the ability to learn new tasks by inferring high-level concepts from existing solutions, then manipulating these concepts in lieu of the raw data. Can we automate this process by deriving latent semantic structures in a document collection using foundation models? We introduce fPLSA, a foundation-model-based Probabilistic Latent Semantic Analysis (PLSA) method that iteratively clusters and tags document segments based on document-level contexts. These tags can be used to model the structure of given documents and for hierarchical sampling of new texts. Our experiments on story writing, math, and multi-step reasoning datasets demonstrate that fPLSA tags help reconstruct the original texts better than existing tagging methods. Moreover, when used for hierarchical sampling, fPLSA produces more diverse outputs with a higher likelihood of hitting the correct answer than direct sampling and hierarchical sampling with existing tagging methods.
@@ 1 Introduction
Large language models (LLMs) have shown impressive performance on a wide range of tasks, such as reasoning (Suzgun et al., 2022; Liu et al., 2023), math problem solving (Wu et al., 2023), and open-ended text generation tasks (Katz et al., 2024; Dubey et al., 2024; OpenAI et al., 2024). Given natural language instructions or in-context examples with chain-of-thought steps, LLMs can adapt quickly to a new task and achieve outstanding performance on challenging tasks that require multi-step reasoning or planning (Wei et al., 2022). However, such methods typically rely on humans to provide the LLM with instructions or chain-of-thought recipes for solving a task. By contrast, humans can directly derive effective methodologies for solving a task by analyzing a separate set of problems and their solutions.
Can we automate the process of discovering latent semantic structures in a document collection using LLMs? Such algorithms would have a wide range of applications, including producing effective guidelines for new tasks, hierarchical sampling for diverse outputs, and document analysis. For example, they can help determine how two document collections differ in text structure and identify the most common plot elements in a story collection.
We frame this problem as an unsupervised clustering and tagging problem, where we discover the text segments that share common characteristics and assign them to the same tag. Based on these segment tags, we can model the latent structure of a collection of documents by learning a dynamic model over the latent tags and their transitions in the documents. Traditional document labeling and topic modeling approaches focus primarily on lexical features such as word or term co-occurrence (Hearst, 1997; Blei et al., 2003; Hofmann et al., 1999), which provide minimal information on the semantics of short text spans. Recent LLM-based approaches discover topics based on higher-level semantic contexts, but rely on one-shot topic generation and merging (Pham et al., 2024; Wang et al., 2023; Mu et al., 2024), which limits the model’s ability to uncover shared characteristics among seemingly unrelated text spans.
@@
In this paper, we introduce fPLSA, an iterative algorithm that alternatively clusters and tags document segments using LLMs based on segment- and document-level contexts. fPLSA combines the merits of traditional topic modeling approaches such as Probabilistic Latent Semantic Analysis (PLSA) (Hofmann et al., 1999) and LLM-based approaches, and captures shared semantic features among text segments more effectively.
We evaluate the informativeness of fPLSA tags by measuring 1) how well they help reconstruct the original text spans, and 2) how useful they are in hierarchical sampling to produce structurally diverse outputs that cover more solution paths. Experiments on story writing, math and multi-step reasoning datasets show that fPLSA leads to higher reconstruction likelihood than existing tagging approaches. Furthermore, on math and reasoning tasks, hierarchical sampling using fPLSA tags produces more diverse outputs, which increase the probability of hitting the correct answer over hierarchical sampling with other tagging approaches.
@@ 2.1 Document Segmentation and Labeling
To model the structure and topic shifts in a document, prior work has introduced unsupervised document segmentation and labeling approaches that leverage term co-occurrence features (Hearst, 1997), co-occurrence shifts in topic vectors (Riedl & Biemann, 2012), lexical features and word embeddings (Glavaš et al., 2016). These approaches focus mostly on lexical features which are limited in modeling the high-level semantic structure of documents. On the other hand, Neural-based approaches have the potential of modeling sentence-level semantics and document-level topic flows more effective, but rely heavily on supervised training samples in the target domain (Koshorek et al., 2018; Arnold et al., 2019; Zhang et al., 2019). Our algorithm infers the structure of documents based on segment- and document-level contexts using LLMs in an unsupervised fashion.
@@ 2.2 Topic Modeling
Topic modeling is a widely used technique in natural language processing for uncovering hidden thematic structures in large text corpora. The most foundational methods in this domain include Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and Probabilistic Latent Semantic Analysis (PLSA) (Hofmann et al., 1999; Hofmann, 1999; 2001). Both methods represent each document as a bag of words and models word-document relationships using a mixture of latent topics, where each topic is represented by a list of top words. These algorithms are mathematically grounded, but typically rely on manual topic interpretation, which often leads to incorrect or incomplete labels (Gillings & Hardie, 2022). More recent work introduces neural topic models (Miao et al., 2016; Dieng et al., 2020; Srivastava & Sutton, 2017), which combine traditional topic models with word embeddings. These models have shown improved performance in handling large and complex vocabularies. However, they still model each document as a bag of words, disregarding the sentence- and document-level semantics. Additionally, the resulting topics are represented either by semantic vectors or lists of closest words, which still rely on manual interpretation. Furthermore, studies have shown that incorporating expert knowledge in topic modeling improves over traditional unsupervised methods (Lee et al., 2017).
@@
However, the generate-and-merge approach limits the model’s potential for discovering shared features among various text spans across documents of different themes and often leads to overly abstract, thematical topics, especially on a large-scale document collection. We propose fPLSA, which combines the merits of traditional PLSA, which uses an iterative EM algorithm to model topic and text distributions, and LLM-based approaches.
@@ 3 Approach
We propose fPLSA, a foundation-model-based EM algorithm that learns the latent tags on a set of segmented documents. We draw inspiration from the traditional Probabilistic Latent Semantic Analysis and use iterative EM steps to learn the latent topics that maximize the estimated likelihood of segmented documents.
@@ 3.1 Probabilistic Latent Semantic Analysis (PLSA)
PLSA models the distribution over words $w$ in a document $d$ as a mixture of conditionally independent multinomial distributions, each such distribution representing a topic $t$. More formally, the generative model of words in a document can be written as:
$p_{\Theta}(w,d)=p(d)\sum_{t}p_{\Theta}(t|d)p_{\Theta}(w|t)$ (1)
where the topic $t$ can be viewed as a discrete latent variable and the total number of discrete topics is pre-defined. $\Theta$ represents the parameters of the PLSA model.
To estimate the parametric distributions $p_{\Theta}(t|d)$ and $p_{\Theta}(w|t)$, PLSA relies on an EM algorithm, which is an iterative method to find the maximum likelihood estimate of parameters in statistical models. Specifically, an EM iteration alternates between an expectation (E) step and a maximization (M) step. At iteration $i$, the E-step estimates the posterior distribution of topics $t$ conditioned on each document $d$ and word $w$ in it based on fixed parameters $\Theta_{i-1}$ from the previous iteration:
@@
The M-step optimizes the parameters $\Theta$ such that the expectation of the joint distribution $p_{\Theta}(w,d)$ with $t$ sampled from the estimated posterior $p_{\Theta_{i-1}}(t|w,d)$ is maximized:
$\Theta_{i}=\arg\max_{\Theta}\mathbb{E}_{t\sim p_{\Theta_{i-1}}(t|w,d)}p(d)p_{\Theta}(t|d)p_{\Theta}(w|t)$ (3)
Theoretically, each EM iteration will yield a larger likelihood in Eq 1 until it converges to a local maximum.
3.2 Foundation-Model-Based PLSA (fPLSA)
@@
We introduce fPLSA, which learns the latent tags (similar to topics in LSA) [^1] We use the terminology tag instead of topic in our algorithm because they may cover shared characteristics among document segments beyond topics (see the example tags in Section 5 for more details). on a set of segmented documents $d=(x_{1},x_{2},...,x_{L})$, where the document $d$ is segmented into $L$ segments $x_{k}$. A core difference between fPLSA and PLSA is that fPLSA models the probability of the sequence of words $(w_{1},w_{2},...,w_{n})$ in each text segment $x_{k}$ jointly as $p_{\Theta}(w_{1},w_{2},...,w_{n}|t)$. Moreover, fPLSA models the distribution over tags $t$ conditioned not only on current segment $x_{k}$ but also on the document $d$. Formally, in fPLSA, the generative model of a segment $x_{k}=w_{1...n}$ in a document $d$ can be written as:
$p_{\Theta}(w_{1...n},x_{k},d)=p(d)p(x_{k}|d)\sum_{t}p_{\Theta}(t|x_{k},d)p_{\Theta}(w_{1...n}|t)$ (4)
@@
Another core difference between fPLSA and PLSA is that we model the parametric distributions $p_{\Theta}(t|x_{k},d)$ and $p_{\Theta}(w_{1...n}|t)$ using an LLM. Specifically, the parameters $\Theta$ in fPLSA include the LLM parameters, which are frozen, and the textual description $\theta_{t}$ for each tag $t$.
Inspired by PLSA, we also maximize the likelihood in Eq 4 using iterative EM steps.
@@
At the E-step in iteration $i$, we approximate the posterior distribution $p_{\Theta_{i-1}}(t|w_{1...n},x_{k},d)$ of tags $t$ conditioned on each document $d$ and segment $x_{k}=w_{1...n}$ in it by prompting the LLM to greedily assign a tag given the tag descriptions $\theta_{i-1}$ from the previous iteration, the current segment $x_{k}=w_{1...n}$ and neighbouring segments $(x_{k-W/2},x_{k+1-W/2},...,x_{k-1+W/2},x_{k+W/2})$ as document-level context, where $W$ is the context window size.
@@
At the M-step, we optimize the tag description $\theta_{t}$ for each tag $t$ by aggregating the segments assigned to tag $t$ and prompting the LLM to generate a tag description that best summarizes what these segments share in common.
@@ 4 Experimental Setup
4.1 Evaluation Datasets
We evaluate fPLSA against various baselines on story writing, math problem solving and multi-step reasoning benchmarks. We use WritingPrompts (Fan et al., 2018), a story writing dataset that contains 300K human-written stories paired with writing prompts from an online forum. We randomly sample 100 stories from the training set for clustering and tagging. We set the number of tags to 100 for all tagging approaches. For math problem solving, we use MATH (Hendrycks et al., 2021), a popular math benchmark that contains high school math competition problems on seven subjects including Prealgebra, Algebra, Number Theory, Counting and Probability, Geometry, Intermediate Algebra and Precalculus. We learn 100 tags on 1K randomly sampled problems and the step-by-step solutions from the training set. We also experiment on the Big-Bench Hard (BBH) benchmark (Suzgun et al., 2022). The original benchmark includes 23 challenging multi-step reasoning tasks, but each task only includes three step-by-step solution examples. Instead, we pick the 12 tasks used in Xu et al. (2024) and use the step-by-step solutions produced by their automatic Chain-of-Thought prompt inference algorithm for clustering and tagging. We set the number of tags to 50 on BBH.
@@ 4.2 Evaluation Metrics
To test how well the learned tags help predict the original texts, we measure the reconstruction log-likelihood of the test documents conditioning on the tags.
Specifically, for each test case $x_{k}$, which is a segment randomly sampled from a test document $x_{1...L}$ (randomly sampled from the test corpus), we measure the reconstruction log-likelihood of $x_{k}$ given latent tags $t_{k}$ under the LLM:
$\mathbb{E}_{t_{k}\sim p_{LLM}(t|x_{1...k-1},x_{k})}[\log p_{LLM}(x_{k}|x_{1...k-1},t_{k})]$ (5)
Specifically, we first sample $S$ alternative segments at position $k$ independently by $\{\hat{x}_{k}^{(1)},\hat{x}_{k}^{(2)},...,\hat{x}_{k}^{(S)}\}\sim p_{LLM}(\cdot|x_{1...k-1})$. Next, we conduct $T$ repeated experiments to approximate the log-likelihood of $x_{k}$ given the previous segments $x_{1…k-1}$ under the LLM. Each time, we randomly sample $C$ alternative segments from $\{\tilde{x}_{k}^{(1)},\tilde{x}_{k}^{(2)},...,\tilde{x}_{k}^{(S)}\}$ and put it together with $x_{k}$ (in randomly shuffled order) as options and ask the LLM which one is the true continuation conditioned on $x_{1...k-1}$ and the tag $t_{k}$ predicted on $x_{k}$. Based on the number of times (denoted as $c_{k}$) that the LLM chooses $x_{k}$ as the true continuation among all $T$ experiments, we estimate the reconstruction log-likelihood with alpha-smoothing ($\alpha=0.1$):
$\mathbb{E}_{t_{k}\sim p_{LLM}(t|x_{1...k-1},x_{k})}[\log p_{LLM}(x_{k}|x_{1...k-1},t_{k})]=\log\frac{c_{k}+\alpha}{T+\alpha S}$ (6)
As a baseline, we compare the reconstruction log-likelihood with the log-likelihood without conditioning on any tags:
$\mathbb{E}[\log p_{LLM}(x_{k}|x_{1...k-1})]$ (7)
which we estimate in the same way as the reconstruction log-likelihood except that when asking the LLM to choose the true continuation, we only provide the previous text segments $x_{1...k-1}$ without any tags.
In our experiments, we estimate the log-likelihood on the same set of 1K randomly sampled test cases using each sampling method.
@@
The latent tags can also be used for hierarchical generation where we first sample a sequence of tags as an outline and then sample the actual text based on the outline. To evaluate if the latent tags help generate more diverse texts, we evaluate if the outputs cover more solution paths and thus lead to higher chance of hitting the correct path on problem solving tasks.
To this end, we evaluate the Hits@K accuracy of hierarchical sampling with latent tags, and compare it with the Hits@K accuracy of direct sampling without tags. Specifically, for each problem, we sample $K=50$ solutions independently from an LLM given the problem description either directly or through hierarchical sampling with latent tags. If any of the $K$ solutions lead to the correct answer, it gets a score of 1, otherwise 0. Finally, we compute the average score over all testing problems.
For hierarchical sampling, we first sample a sequence of tags $(t_{1},t_{2},...,t_{l})$ (up till the special tag <END>) with maximum length L using a bigram model learned on the training data (based on the tag assignments):
$p(t_{1},t_{2},...,t_{l})=p(t_{1})p(t_{2}|t_{1})...p(t_{l}|t_{l-1})p(\text{<END>}|t_{l})$ (8)
And then, we prompt the LLM to sample a solution to the given problem based on the sampled sequence of tags $(t_{1},t_{2},...,t_{l})$.
4.3 fPLSA Setup
For the EM procedure, we set the maximum number of iterations to 30. At the E-step (where the LLM assigns a tag to each segment conditioned not only on the current segment but also on neighbouring segments within the context window), we use a context window size of 2 on WritingPrompts and use unlimited context window (such that the whole solution is used as context) on MATH and BBH. At the M-step, we randomly sample 10 segments assigned to each tag to update the tag description.

### v2

# fLSA: Learning Semantic Structures in Document Collections Using Foundation Models

@@ Abstract
Humans can learn to solve new tasks by inducing high-level strategies from example solutions to similar problems and then adapting these strategies to solve unseen problems. Can we use large language models to induce such high-level structure from example documents or solutions? We introduce fLSA, a foundation-model-based Latent Semantic Analysis method that iteratively clusters and tags document segments based on document-level contexts. These tags can be used to model the latent structure of given documents and for hierarchical sampling of new texts. Our experiments on story writing, math, and multi-step reasoning datasets demonstrate that fLSA tags are more informative in reconstructing the original texts than existing tagging methods. Moreover, when used for hierarchical sampling, fLSA tags help expand the output space in the right directions that lead to correct solutions more often than direct sampling and hierarchical sampling with existing tagging methods.
fLSA: Learning Semantic Structures in Document Collections
Using Foundation Models
Weijia Xu, Nebojsa Jojic & Nicolas Le Roux Microsoft Research Redmond, WA 98052, USA {weijiaxu,jojic,nleroux}@microsoft.com
@@ 1 Introduction
Figure 1: Visualizing the bigram dynamic model over the latent tags learned on MATH solutions. For each tag, we list the three most probable next tags based on the transition probabilities $p(t_{k}|t_{k-1})$. The transition probabilities are annotated on the arrows. For Tag 24, we also list two example next tags outside the top-3 choices with transition probabilities $p\approx 0.01$.
Figure 2: An example of using the sampled tag sequence as an outline (in purple) to aid an LLM in generating a solution (italicized) to the given problem (in blue).
@@
Large language models (LLMs) have shown impressive performance on a wide range of tasks, such as reasoning (Suzgun et al., 2022; Liu et al., 2023), math problem solving (Wu et al., 2023), and open-ended text generation tasks (Katz et al., 2024; Dubey et al., 2024; OpenAI et al., 2024). Given natural language instructions or in-context examples with chain-of-thought steps, LLMs can adapt quickly to a new task and achieve outstanding performance on challenging tasks that require multi-step reasoning or planning (Wei et al., 2022). However, such methods typically rely on humans to induce the common strategy for solving a type of problems and demonstrate the strategy through few-shot chain-of-thought prompting. By contrast, humans learn to solve a new type of problems by analyzing some example problems and their solutions, inducing the common strategies (i.e. latent semantic structure) underlying these problem solutions, and testing them out on the new problems.
Inducing the latent semantic structure in a set of documents can be modeled as an unsupervised clustering and tagging problem, where given a set of coarsely segmented documents, we cluster the text segments that share common characteristics into the same set and assign a tag to each set of segments. Based on these segment tags, we can then uncover the latent structure by learning a dynamic model over the latent tags and their transition probabilities in the document set. As an example, Figure 1 shows a dynamic model over learned tags in mathematical solutions. Such dynamic models can help humans better understand and analyze large collections of documents. They also encode more generalizable information compared to few-shot examples, providing a useful guide for LLMs to solve new problems without manual intervention (as shown by the example in Figure 2). Additionally, they can also aid in searching algorithms on complex reasoning tasks (Guan et al., 2025) through hierarchical sampling: one can sample from the dynamic model over latent tags as an outline for the actual solution steps to explore more diverse solution paths during the rollout stage.
@@
In this paper, we introduce fLSA, an iterative algorithm that alternatively clusters and tags document segments using LLMs based on segment- and document-level contexts. fLSA combines the merits of traditional topic modeling approaches such as Latent Semantic Analysis (LSA) (Hofmann et al., 1999) and LLM-based approaches, and captures shared semantic features among text segments more effectively. We evaluate 1) the informativeness of fLSA tags by measuring how well they help reconstruct the original text spans, and 2) their usefulness in expanding the search space in the right directions by measuring the Hits@K accuracy of the generated solutions through hierarchical sampling using the tags. Experiments on story writing, math and multi-step reasoning datasets show that fLSA leads to higher reconstruction likelihood than existing tagging approaches. Furthermore, on math and reasoning tasks, hierarchical sampling using fLSA tags helps expand the output space in the right directions more effectively than both direct sampling and existing tagging methods.
@@ 2.1 Document Segmentation and Labeling
To model the structure and topic shifts in a document, prior work has introduced unsupervised document segmentation and labeling approaches that leverage term co-occurrence features (Hearst, 1997), co-occurrence shifts in topic vectors (Riedl and Biemann, 2012), lexical features and word embeddings (Glavaš et al., 2016). These approaches focus mostly on lexical features which are limited in modeling the high-level semantic structure of documents. On the other hand, Neural-based approaches have the potential of modeling sentence-level semantics and document-level topic flows more effective, but rely heavily on supervised training samples in the target domain (Koshorek et al., 2018; Arnold et al., 2019; Zhang et al., 2019). Our algorithm infers the structure of documents based on segment- and document-level contexts using LLMs in an unsupervised fashion.
@@ 2.2 Topic Modeling
Topic modeling is a widely used technique in natural language processing for uncovering hidden thematic structures in large text corpora. The most foundational methods in this domain include Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and Latent Semantic Analysis (LSA) (Hofmann et al., 1999; Hofmann, 1999, 2001). Both methods represent each document as a bag of words and models word-document relationships using a mixture of latent topics, where each topic is represented by a list of top words. These algorithms are mathematically grounded, but typically rely on manual topic interpretation, which often leads to incorrect or incomplete labels (Gillings and Hardie, 2022). More recent work introduces neural topic models (Miao et al., 2016; Dieng et al., 2020; Srivastava and Sutton, 2017), which combine traditional topic models with word embeddings. These models have shown improved performance in handling large and complex vocabularies. However, they still model each document as a bag of words, disregarding the sentence- and document-level semantics. Additionally, the resulting topics are represented either by semantic vectors or lists of closest words, which still rely on manual interpretation. Furthermore, studies have shown that incorporating expert knowledge in topic modeling improves over traditional unsupervised methods (Lee et al., 2017).
@@
However, the generate-and-merge approach limits the model’s potential for discovering shared features among various text spans across documents of different themes and often leads to overly abstract, thematical topics, especially on a large-scale document collection. We propose fLSA, which combines the merits of traditional LSA, which uses an iterative EM algorithm to model topic and text distributions, and LLM-based approaches.
@@ 3 Approach
Figure 3: An illustration of the E-step and M-step in fLSA. At the E-step, we assign each text segment to a tag through prompting given the tag descriptions at the previous iteration. At the M-step, we prompt the LLM to generate new tag descriptions based on the segments assigned to each tag at the E-step.
We propose fLSA, a foundation-model-based EM algorithm that learns the latent tags on a set of segmented documents. We draw inspiration from the traditional Probabilistic Latent Semantic Analysis and use iterative EM steps to learn the latent tags that maximize the estimated likelihood of segmented documents.
@@ 3.1 Probabilistic Latent Semantic Analysis (PLSA)
PLSA models the distribution over words $w$ in a document $d$ as a mixture of conditionally independent multinomial distributions, each such distribution representing a topic $t$. This generative model of words in a document is usually expressed mathematically in terms of the distribution:
$p_{\Theta}(w|d)=\sum_{t}p_{\Theta}(t|d)p_{\Theta}(w|t),$ (1)
which can be sampled by first sampling a topic $t$ for the given document $d$ from $p_{\Theta}(t|d)$ and then sampling words conditioned on the topic from $p_{\Theta}(w|t)$. $\Theta$ represents the parameters of the PLSA model. PLSA aims to find $\Theta$ that maximizes the log-likelihood of words in all documents:
$\mathcal{L}=\sum_{d,w}\log\sum_{t}p_{\Theta}(t|d)p_{\Theta}(w|t)$ (2)
To estimate the parametric distributions $p_{\Theta}(t|d)$ and $p_{\Theta}(w|t)$, PLSA relies on an EM algorithm, which is an iterative method to find the maximum likelihood estimate of parameters in statistical models. Specifically, an EM iteration alternates between an expectation (E) step and a maximization (M) step. At iteration $i$, the E-step estimates the posterior distribution $p_{\Theta_{i-1}}(t|w,d)$ of topics $t$ conditioned on each document $d$ and word $w$ in it based on fixed parameters $\Theta_{i-1}$ from the previous iteration:
@@
The M-step optimizes the parameters $\Theta$ such that the expectation of the log-likelihood $p_{\Theta}(w|d)$ of words in each document given $t$ sampled from the estimated posterior $p_{\Theta_{i-1}}(t|w,d)$ is maximized:
$\arg\max_{\Theta}\sum_{d,w}\mathbb{E}_{t\sim p_{\Theta_{i-1}}(t|w,d)}\log p_{\Theta}(t|d)p_{\Theta}(w|t)$ (4)
Theoretically, each EM iteration will yield a larger likelihood in Eq 2 until it converges to a local maximum. In topic modeling literature, various generalized EM variants exist, including the ones that approximate the posterior distribution with a small number of samples, or just the mode of it, and which alter the parameters so that they do not necessarily maximize the likelihood under the posterior, but simply improve it.
3.2 Foundation-Model-Based LSA (fLSA)
@@
We introduce fLSA, which learns the latent tags (similar to topics in LSA) [^1] We use the terminology tag instead of topic in our algorithm because they may cover shared characteristics among document segments beyond topics (see the example tags in Figure 1). on a set of segmented documents $d=(x_{1},x_{2},...,x_{L})$, where the document $d$ is segmented into $L$ segments $x_{k}$. A core difference between fLSA and PLSA is that PLSA models the generative probability of each word in a document independently, while fLSA models the probability of the sequence of words $(w_{1},w_{2},...,w_{n})$ in each text segment $x_{k}$ jointly as $p_{\Theta}(w_{1},w_{2},...,w_{n}|t)$. Moreover, PLSA models the distribution over tags $p_{\Theta}(t|d)$ for each document independently of other documents, while fLSA models the distribution over tags $t$ conditioned not only on current segment $x_{k}$ but also on the document $d$.
To express the difference mathematically, in fLSA, the generative model of a segment $x_{k}=w_{1..n}$ in a document $d$ can be written as:
$p_{\Theta}(w_{1..n}|x_{k},d)=\sum_{t}p_{\Theta}(t|x_{k},d)p_{\Theta}(w_{1..n}|t),$ (5)
which can be sampled by first sampling a tag $t$ for the current segment $x_{k}$ in document $d$ and then sampling the word sequence $w_{1..n}$ for that segment given the tag.
@@
Another core difference between fLSA and PLSA is that we model the parametric distributions $p_{\Theta}(t|x_{k},d)$ and $p_{\Theta}(w_{1..n}|t)$ using an LLM with frozen parameters, and the tunable “parameters” $\Theta$ in fLSA are the textual description $\Theta(t)$ for each tag $t$ and the tag assignment for each segment.
Analogously to the (generalized) EM algorithms for traditional topic models, we are seeking  $\Theta$ that corresponds to high likelihood of the word sequence in each document:
$\mathcal{L}=\sum_{d,x_{k}}\log\sum_{t}p_{\Theta}(t|x_{k},d)p_{\Theta}(w_{1..n}|t)$ (6)
@@
Our iterative EM steps are shown in Figure 3. At the E-step in iteration $i$, we approximate the posterior distribution $p_{\Theta_{i-1}}(t|w_{1..n},x_{k},d)$ of tags $t$ for each segment $x_{k}=w_{1..n}$ in document $d$ by prompting the LLM to greedily assign a tag given the tag descriptions $\Theta_{i-1}(t)$ from the previous iteration, the current segment $x_{k}=w_{1..n}$ and neighbouring segments $(x_{k-W/2},x_{k+1-W/2},...,x_{k-1+W/2},x_{k+W/2})$ as document-level context, where $W$ is the context window size. [^2] At the first iteration, since the tag descriptions are empty, we assign tags randomly. At the M-step, in lieu of maximizing (or just improving) the expected log-likelihood $p_{\Theta}(w_{1..n}|x_{k},d)$ of words in each segment given the tag assignments from the E-step,
$\begin{split}\arg\max_{\Theta}\sum_{d,x_{k}}&\mathbb{E}_{t\sim p_{\Theta_{i-1}}(t|w_{1..n},x_{k},d)}\\ &\log p_{\Theta}(t|x_{k},d)p_{\Theta}(w_{1..n}|t),\end{split}$ (7)
@@
we obtain updated tag descriptions $\Theta(t)$ by inviting the LLM itself to summarize the segments assigned to the tag $t$: We aggregate the segments assigned to tag $t$ and prompt the LLM to generate a tag description that best summarizes what these segments share in common (Fig. 3).
@@ 4 Experimental Setup
4.1 Datasets
We evaluate fLSA against various baselines on WritingPrompts (a story writing dataset (Fan et al., 2018)), MATH (which contains math problems and the corresponding solution texts (Hendrycks et al., 2021)), and Big-Bench Hard (BBH) benchmark (which contains diverse types of reasoning problems and their solutions (Suzgun et al., 2022)). We set the number of tags to 100 for WritingPrompts and MATH, and 50 for BBH (see the Appendix for more details).
@@ 4.2 Evaluation Metrics
To measure the informativeness of learned tags (either through fLSA or a baseline algorithm), we measure the reconstruction log-likelihood of the test documents (stories in the test set of WritingPrompts or problem solutions in the test set of MATH) conditioned on the tags.
Specifically, for each test case $x_{k}$, which is a segment randomly sampled from a test document $x_{1...L}$ (randomly sampled from the test corpus), we approximate the reconstruction log-likelihood of $x_{k}$ given latent tags $t_{k}$ predicted given $x_{k}$ and its neighboring segments under the LLM:
$\mathbb{E}_{t_{k}\sim p_{LLM}(t|x_{k},d)}[\log p_{LLM}(x_{k}|x_{1...k-1},t_{k})]$ (8)
Specifically, we first sample $S$ alternative segments at position $k$ independently by $\{\tilde{x}_{k}^{(1)},\tilde{x}_{k}^{(2)},...,\tilde{x}_{k}^{(S)}\}\sim p_{LLM}(\cdot|x_{1...k-1})$. Next, we conduct $T$ repeated experiments to approximate the log-likelihood of $x_{k}$ given the previous segments $x_{1…k-1}$ and the tag $t_{k}$ predicted on $x_{k}$ under the LLM. Each time, we randomly sample $C$ alternative segments from $\{\tilde{x}_{k}^{(1)},\tilde{x}_{k}^{(2)},...,\tilde{x}_{k}^{(S)}\}$ and put it together with $x_{k}$ (in randomly shuffled order) as options and ask the LLM which one is the true continuation conditioned on $x_{1...k-1}$ and $t_{k}$. Based on the number of times (denoted as $c_{k}$) that the LLM chooses $x_{k}$ as the true continuation among all $T$ experiments, we estimate the reconstruction log-likelihood with alpha-smoothing ($\alpha=0.1$):
$\begin{split}&\mathbb{E}_{t_{k}\sim p_{LLM}(t|x_{k},d)}[\log p_{LLM}(x_{k}|x_{1...k-1},t_{k})]\\ &=\log\frac{c_{k}+\alpha}{T+\alpha S}\end{split}$ (9)
As a baseline, we compare the reconstruction log-likelihood with the log-likelihood computed the same way as above but without conditioning on any tags:
$\mathbb{E}[\log p_{LLM}(x_{k}|x_{1...k-1})]=\log\frac{c_{k}^{\prime}+\alpha}{T+\alpha S}$ (10)
where $c_{k}^{\prime}$ is the number of times that the LLM chooses $x_{k}$ as the true continuation among $T$ experiments, which is computed the same way as above except that when asking the LLM to choose the true continuation, we only provide the previous text segments $x_{1...k-1}$ without any tags.
In our experiments, we evaluate the reconstruction log-likelihood of all methods on the same set of 1K randomly sampled test cases.
@@
To demonstrate that the learned tags can also help expand the search space in the right directions when searching for effective solutions to a complex reasoning task, we learn a dynamic model over the latent tags (as shown by the example in Figure 1) and use it for hierarchical sampling, where we first sample a sequence of tags as an outline and then sample the actual text based on the outline. And then, we evaluate the Hits@K accuracy of hierarchical sampling with latent tags, and compare it with the Hits@K accuracy of direct sampling without tags. Specifically, for each problem, we sample $K=50$ solutions independently from an LLM given the problem description either directly or through hierarchical sampling with latent tags. If any of the $K$ solutions leads to the correct answer, it gets a score of 1, otherwise 0. Finally, we compute the average score over all testing problems.
For hierarchical sampling, we first sample a sequence of tags $(t_{1},t_{2},...,t_{l})$ (up till the special tag <END>) with maximum length L using a bigram model learned on the training data (without conditioning on the test problem):
$\begin{split}&p(t_{1},t_{2},...,t_{l})\\ =&p(t_{1})p(t_{2}|t_{1})...p(t_{l}|t_{l-1})p(\text{<END>}|t_{l})\end{split}$ (11)
And then, we prompt the LLM to generate a solution to the given problem based on the tag sequence $(t_{1},t_{2},...,t_{l})$ using the prompt template shown in Figure 2.
4.3 fLSA Setup
For the EM procedure, we set the maximum number of iterations to 30.333We found in our preliminary experiment that the learned tag descriptions become stable (with very little semantic changes) in less than 30 iterations. At the E-step (where the LLM assigns a tag to each segment conditioned not only on the current segment but also on neighbouring segments within the context window), we use a context window size of 2 on WritingPrompts and use unlimited context window (such that the whole solution is used as context) on MATH and BBH. At the M-step, we randomly sample 10 segments assigned to each tag to update the tag description.